OpenStudy (anonymous):

Evaluate The Equation : 15 d = -105 Can someone tell me what this is?

OpenStudy (the_fizicx99):

15d = -105 To isolate D, do the inverse ( Multiplication = division) So divide both sides by 15, 15d = -105 /15 /15 d = ??

OpenStudy (anonymous):

thank you :)

OpenStudy (the_fizicx99):

Np, did you figure it out?

OpenStudy (anonymous):

ok i did 15 divided by 15 and i got 1 then i did 105 divided by 15 and i got 7 what do i do nexted

OpenStudy (the_fizicx99):

That's it, d = -7 15 * (-7) = -105

OpenStudy (anonymous):

ohhh thank you

OpenStudy (the_fizicx99):

np :p